OpenStudy (anonymous):

What are the coordinates of the vertex of the function below? Write your answer in the form (x,y). y+9=-6(x-2)^2

OpenStudy (anonymous):

missing some information are you not?

OpenStudy (anonymous):

y+9=-6(x^2-4x+4)

OpenStudy (anonymous):

y+9= -6x^2-24x+-24

OpenStudy (anonymous):

y=-6x^2-24x-33

OpenStudy (anonymous):

nope thats the whole question

OpenStudy (anonymous):

ok then set the line x,y to 0

OpenStudy (anonymous):

0=-6x^2-24x-33

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

oops equation is wrong sorry

OpenStudy (anonymous):

-6x^2+24x-33

OpenStudy (anonymous):

derivative?

OpenStudy (anonymous):

what about the derivative?

OpenStudy (anonymous):

vertex is \[(2,-9)\] from your eyeballs

OpenStudy (anonymous):

don't multiply out. don't take the derivative. just look

OpenStudy (anonymous):

lol, but that's the easy way!! we have all these shiny tools to use!! lol

OpenStudy (anonymous):

you gotta be joking. all that work basically is to put it in the format \[y=-6(x-2)^2-9\] so you can read off the answer. you are one step away